A has mass 40 kg, B is of 15 kg and F is 500 N parallel to a smooth inclined plane. The system is moving together. Show that the acc of the system is 45/11 m /s2  and the least coeff  of friction  b/w A and B is 9 (3)1/2  / 53.

Grade:12

1 Answers

Ramesh V
14 years ago

Free body diagrams of two blocks are shown below:

on balancing forces on each block in X and y axis respectively, we have:

for A:

N -150 = 7.5a

f = uN =13a

for B:

0.866R - N =150 + 20a

433 -0.5R -uN = 34.64a   here u is coeff. of friction btn 2 blocks

on solving above equations, we have

N = 150 + 7.5a

433 - 47.64a = 0.5R

so, its     750 - 82.5a = 300 + 27.5a

450 = 110a

the acc of the system a = 45 / 11 m/sec2

the coeff. of friction is

u(150 + 20a ) = 13a

so u = 13a / (150+20a)

The coeff. of friction is    u = 0.2942 or 9.(3)1/2 / 53.

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch